# Gravitational wave inference of star cluster properties from intermediate-mass black hole mergers

### v1

@@ I Introduction
While the current LIGO-Virgo-KAGRA network of gravitational wave (GW) observatories routinely detects compact binary coalescences Abbott et al. (2019, 2021a, 2021b), the community is developing next-generation (XG) ground-based GW observatories, such as the Einstein Telescope (ET) Punturo et al. (2010) and Cosmic Explorer (CE) Evans et al. (2021). With improved sensitivity at frequencies of the order of a few Hz, these detectors will observe the mergers of black holes (BHs) with masses in the hundreds of solar masses out to high redshift Branchesi et al. (2023a); Gupta et al. (2023); Fairhurst et al. (2023); Reali et al. (2024). As such, these detectors will give us access to the lower end of the intermediate-mass black hole (IMBH) regime, bridging the gap between stellar mass and supermassive BHs. Understanding the IMBH population and their environments is important, because IMBHs are believed to seed the growth of supermassive BHs Volonteri (2012).
@@
Several mechanisms can lead to the formation and merger of IMBH-IMBH binaries. One of the scenarios for BH growth is the hierarchical assembly of BHs, first from the stellar mass to the intermediate mass regime, and then to the supermassive regime Ebisuzaki et al. (2001); Sedda and Mastrobuono-Battisti (2019); Askar et al. (2021). In this scenario, IMBHs first form from gravitational runaways in the cores of dense star clusters. Clusters that reside within a few kpcs from the center of their host galaxy undergo dynamical friction and can sink into the core before they ”evaporate” (Antonini, 2013), where they merge with the nuclear star cluster or with other inspiralling globular clusters (see e.g. Sec. 7.2 of Ref. Neumayer et al. (2020)). Dynamical friction also contributes to the growth of the central nuclear star cluster Antonini (2013); Gnedin et al. (2014). In this way, IMBHs assembled in different cluster environments are brought together as their host clusters carry them along, eventually ending up in the central regions of the galaxy Arca-Sedda and Gualandris (2018). For example, the compact stellar complex IRS 13E is hypothesized to be the core of a dissolved young star cluster hosting an IMBH that sank into the Galactic center. However, the absence of a radio or X-ray signature casts doubt on the presence of an IMBH at the center of this system Schödel et al. (2005), that could be dominated by a subcluster of stellar-mass BHs Banerjee and Kroupa (2011). A second possible scenario leading to the formation of IMBH-IMBH pairs is the merger of ultra-dwarf galaxies. Dynamical friction brings the IMBHs into the central regions, and the sinking timescale can be significantly shorter if the IMBHs are embedded in NSCs Palmese and Conselice (2021). According to observations in the local Universe, the NSC occupation fraction (i.e., the fraction of systems with identified NSCs) is at least 80% in the galaxy stellar mass range $\sim 10^{8}$$10^{10}M_{\odot}$: see the right panel of Fig. 3 in Neumayer et al. (2020); Rantala et al. (2024). In a third scenario, star clusters can form in the clumpy star-forming turbulent environment of giant molecular cloud complexes, and then merge hierarchically after sinking into the central regions. This process can be very efficient and rapid, as suggested by magnetohydrodynamic simulations Shi et al. (2024). In particular, IMBHs can be brought together and form a binary much more efficiently when surrounded by stellar clusters, as the dynamical friction timescale can be effectively reduced by orders of magnitude for the cluster-IMBH system due to its higher dynamical mass (see also Fig. 3 of Palmese and Conselice (2021)). The simulations of Ref. Khan and Holley-Bockelmann (2021) suggest that the formation and hardening of IMBH-IMBH pairs in the centers of dwarfs is very efficient and leads to their merger.
@@ II.1 Stellar consumption rate
Assuming most of the cluster’s mass is locked in light stars that survive for billions of years, the stellar number is $N_{\star}\simeq M_{\rm cl}/\overline{m}_{\star}$ where $\overline{m}_{\star}\simeq 0.6M_{\odot}$ for a Kroupa initial mass function Kroupa (2002). We neglect the contribution of BHs (and of the central BH, of mass $M_{\rm BH}$) to the total mass of the cluster, as the former can be shown to contribute no more than $\simeq 10\%$ of the total mass Kritos and Cholis (2020). According to the virial theorem, the root-mean-square velocity in the cluster is $v_{\rm rms}\simeq(0.4GN_{\star}\overline{m}_{\star}/r_{\rm h})^{1/2}$ [see e.g. Eq. (2) from Ref. Spitzer and Hart (1971)]. We further assume isothermal and isotropic conditions within the half-mass radius, so that the velocity dispersion in the cluster core is $\sigma=v_{\rm rms}/\sqrt{3}$. The influence radius of the central BH is defined as $r_{\rm a}\equiv GM_{\rm BH}/\sigma^{2}$, and it corresponds approximately to the radius within which the potential of the BH starts to dominate over the stellar component. It also roughly corresponds to the distance from the center which encloses a total stellar mass of $\approx M_{\rm BH}$. Stars within $r_{\rm a}$ can have bound orbits around the central BH.
@@ II.2 Star cluster evolution
The tidally limited model can be applied to systems that experience the tidal field of the host galaxy within the inner kpc and inspiral to the center. On the other hand, the isolated model is more appropriate for the evolution of nuclear star clusters, since the tidal field cancels at the center of the galaxy. The isolated cluster evolution model is also adequate to describe clusters that reside several kpc away from the galaxy center; however, these isolated clusters in the halo of the galaxy are unlikely to merge with other clusters due to small coalescence cross-sections.
@@
which can be integrated to get the time evolution of the stellar number:
@@
Substituting (6) into (4) then yields $r_{\rm h}(t)$. When $\xi_{\rm e}=0$ and $\nu=0$ (i.e., in the case of no mass loss and no stellar evolution) this solution reduces to the late-time radius growth law $r_{\rm h}(t)\propto(t-\tau_{\rm cc})^{2/3}$: see e.g. Eq. (4) in Antonini et al. (2019).
@@ II.4 Dependence on the cluster’s initial conditions
Figure 3: Contour plots of the BH mass assembled via runaway tidal encounters in the $(M_{\rm cl,0},r_{\rm h,0})$ plane at $t=0.1\tau_{\rm Hub}$ (left panels) and $t=\tau_{\rm Hub}$ (right panels). The top and bottom panels refer to isolated and tidally limited star clusters, respectively. The $\tau_{\rm BH}\geq\tau_{\rm Hub}$ ($\tau_{\rm BH}\geq 0.1\,\tau_{\rm Hub}$) region above the blue dashed line corresponds to clusters for which the time required for the BH to begin growing is larger than a Hubble time (one-tenth of a Hubble time) due to non-evaporation of the BH subsystem. In the top panels, the $\tau_{\rm cc}\leq 2\,\rm Myr$ region below the dashed orange line corresponds to clusters for which core collapse occurs before massive stars become BHs, and an IMBH might form through runaway stellar mergers instead. The $v_{\rm esc,0}>200\,\rm km\,s^{-1}$ region (dashed green line) corresponds to clusters for which the initial escape velocity is large enough for the repeated BH merger channel to dominate. In the bottom panels, the $\tau_{\rm ev}\leq\tau_{\rm Hub}$ ($\tau_{\rm ev}\leq 0.1\tau_{\rm Hub}$) region corresponds to clusters that evaporate on a timescale less than $\tau_{\rm Hub}$ ($0.1\tau_{\rm Hub}$): see Eq. (7). The parameter $f_{\rm s}$ is set to 0.5.
@@ II.5 Delay time
We consider the formation of IMBH-IMBH pairs through the coalescence of two star clusters, each cluster hosting one IMBH (see Fig. 1). This can be accomplished either through the merger of globular clusters that inspiral into the central regions of their galaxy through dynamical friction Ebisuzaki et al. (2001) or through the coalescence of nuclear star clusters in the post-merger phase of two dwarf galaxies Palmese and Conselice (2021). Alternatively, IMBHs may form in star clusters that fragment in the turbulent central regions of a galaxy and are then brought together to form a bound pair Shi et al. (2024).
@@
    If $E_{\rm b}=Gm_{1}m_{2}/(2a)$ is the binding energy of the IMBH-IMBH binary, then the rate of change of binding energy is given by $dE_{\rm b}/dt=\langle\Delta E_{\rm b}\rangle(dN/dt)$. Here the mean energy change per interaction is given by $\langle\Delta E_{\rm b}\rangle=(H/(4\pi))(m_{\star}/m_{12})E_{\rm b}$, where $H\simeq 15$ and $m_{12}=m_{1}+m_{2}$ Quinlan (1996). Moreover, $dN/dt$ is the binary-single interaction rate given by Eq. (6.15) of Ref. Merritt and Poon (2004), computed by setting the loss-cone radius equal to the semimajor axis of the binary, $a_{\rm B}$, and the central mass equal to the binary’s mass, $m_{12}$. Combining these equations leads to a differential equation for the evolution of $a_{\rm B}$:
@@
    The semi-major axis shrinks until the emission of GWs is efficient enough to drive the evolution of the binary. At this point, the binary decouples from the dynamics, and we denote the corresponding semimajor axis by $a_{\rm gw}$. The eccentricity at decoupling, $e_{\rm gw}$, is sampled from a thermal distribution, and the hardening timescale $\tau_{\rm har}$ is computed as in Eq. (9) from Ref. Sesana and Khan (2015), where the density is set equal to $\rho_{\rm eff}$.
@@ III.1 Bayesian analysis
In this subsection we present the Bayesian framework used to infer the cluster properties. We start with a description of the cluster parameters (Sec. III.1.1), followed by the method we use to reweigh the prior (Sec. III.1.2)) and a discussion of our chosen astrophysical prior (Sec. III.1.3).
@@ III.1.2 Source-parameter prior reweighing
When performing PE on our chosen events, we assume some prior on the source parameters, denoted by $p_{\rm PE}(\boldsymbol{\theta})$. Since the likelihood function remains invariant under the $\cal F$-map, so that $p(d|\boldsymbol{\theta})=p(d|\boldsymbol{\lambda})$, it follows that the priors on $\boldsymbol{\theta}$ and $\boldsymbol{\lambda}$ are related by a Jacobian as $p_{\rm PE}(\boldsymbol{\theta})d\boldsymbol{\theta}=p_{\rm PE}(\boldsymbol{\lambda})d\boldsymbol{\lambda}$, and therefore they are not independent from each other. This implies that a prior on $\boldsymbol{\lambda}$ is induced by the PE prior choice, which may not be consistent with the astrophysically motivated prior $p_{\rm astro}(\boldsymbol{\lambda})$ (that will be described in Sec. III.1.3 below). Hence, we reweigh the posterior samples on source parameters of our merger events by the astrophysical prior on $\boldsymbol{\theta}$, $p_{\rm astro}(\boldsymbol{\theta})$, which we introduce to be consistent with $p_{\rm astro}(\boldsymbol{\lambda})$. To obtain $p_{\rm astro}(\boldsymbol{\theta})$, we first draw samples $\{\boldsymbol{\lambda}_{i}\}$ from $p_{\rm astro}(\boldsymbol{\lambda})$, then we get the samples $\boldsymbol{\theta}_{i}={\cal F}(\boldsymbol{\lambda}_{i})$ with forward modelling, and finally we estimate the function $p_{\rm astro}(\boldsymbol{\theta})$ with kernel density estimation on these samples.
@@ IV.1.2 Runaway stellar collisions
A simple analytic model for the production of IMBHs through this channel, where the mass of the IMBH progenitor is given as a function of the cluster’s initial mass, was developed in Ref. Portegies Zwart and McMillan (2002) and verified through $N$-body simulations. Assuming that the very massive star assembled via this mechanism collapses directly into an IMBH, and neglecting any stellar wind effects that would reduce the mass of the star by the end of its life, Eq. (16) from Portegies Zwart and McMillan (2002) finds the simple relation $M_{\rm IMBH}\simeq m_{\rm seed}+0.008M_{\rm cl,0}$. Here $m_{\rm seed}$ is the zero-age main-sequence seed mass of the heaviest star that initially forms in the cluster, which may be $\sim 150M_{\odot}$ Figer (2005). Based on this equation, the mass scale of star clusters that contribute to the formation of hundred-$M_{\odot}$ IMBHs is similar to the mass scale of the runaway TDE channel examined in this paper for $f_{\rm s}=0.5$. Nevertheless, the stellar collision channel requires sufficiently small initial relaxation times (hence, very compact stellar systems), so that the process can kick-start before the heaviest stars form a BH subsystem that can prevent such a stellar runaway. The formation of the IMBH through stellar collisions is very rapid (within only a few Myr), and based on the classification of Ref. Giersz et al. (2015) it would correspond to a “fast” growth scenario.
@@ IV.1.4 Gas accretion
Observational data imply that, despite being dry stellar systems, globular clusters contain some gas in their centers, resulting from stellar winds of low-mass stars and ionized by ambient white dwarfs. For instance, Reference Abbate et al. (2018) inferred a density of gas particles $n_{\rm e}=0.23\,\rm cm^{-3}$ using the dissipative effects that the gas would have on the emission of millisecond pulsars in 47 Tucanae.

### v2

@@ I Introduction
While the current LIGO-Virgo-KAGRA network of gravitational wave (GW) observatories routinely detects compact binary coalescences Abbott et al. (2019, 2021, 2023), the community is developing next-generation (XG) ground-based GW observatories, such as the Einstein Telescope (ET) Punturo et al. (2010) and Cosmic Explorer (CE) Evans et al. (2021). With improved sensitivity at frequencies of the order of a few Hz, these detectors will observe the mergers of black holes (BHs) with masses in the hundreds of solar masses out to high redshift Branchesi et al. (2023a); Gupta et al. (2024); Fairhurst et al. (2024); Reali et al. (2024). As such, these detectors will give us access to the lower end of the intermediate-mass black hole (IMBH) regime, bridging the gap between stellar mass and supermassive BHs. Understanding the IMBH population and their environments is important, because IMBHs are believed to seed the growth of supermassive BHs Volonteri (2012).
@@
Several mechanisms can lead to the formation and merger of IMBH-IMBH binaries. One of the scenarios for BH growth is the hierarchical assembly of BHs, first from the stellar mass to the intermediate mass regime, and then to the supermassive regime Ebisuzaki et al. (2001); Sedda and Mastrobuono-Battisti (2019); Askar et al. (2021). In this scenario, IMBHs first form from gravitational runaways in the cores of dense star clusters. Clusters that reside within a few kpcs from the center of their host galaxy undergo dynamical friction and can sink into the core before they “evaporate” (Antonini, 2013), where they merge with the nuclear star cluster or with other inspiralling globular clusters (see, e.g., Sec. 7.2 of Ref. Neumayer et al. (2020)). Dynamical friction also contributes to the growth of the central nuclear star cluster Antonini (2013); Gnedin et al. (2014). In this way, IMBHs assembled in different cluster environments are brought together as their host clusters carry them along, eventually ending up in the central regions of the galaxy Arca-Sedda and Gualandris (2018). For example, the compact stellar complex IRS 13E is hypothesized to be the core of a dissolved young star cluster hosting an IMBH that sank into the Galactic center. However, the absence of a radio or x-ray signature casts doubt on the presence of an IMBH at the center of this system Schödel et al. (2005), that could be dominated by a subcluster of stellar-mass BHs Banerjee and Kroupa (2011). A second possible scenario leading to the formation of IMBH-IMBH pairs is the merger of ultradwarf galaxies. Dynamical friction brings the IMBHs into the central regions, and the sinking timescale can be significantly shorter if the IMBHs are embedded in NSCs Palmese and Conselice (2021). According to observations in the local Universe, the NSC occupation fraction (i.e., the fraction of systems with identified NSCs) is at least 80% in the galaxy stellar mass range $\sim 10^{8}$$10^{10}M_{\odot}$: see the right panel of Fig. 3 in Neumayer et al. (2020); Rantala et al. (2024). In a third scenario, star clusters can form in the clumpy star-forming turbulent environment of giant molecular cloud complexes, and then merge hierarchically after sinking into the central regions. This process can be very efficient and rapid, as suggested by magnetohydrodynamic simulations Shi et al. (2024). In particular, IMBHs can be brought together and form a binary much more efficiently when surrounded by stellar clusters, as the dynamical friction timescale can be effectively reduced by orders of magnitude for the cluster-IMBH system due to its higher dynamical mass (see also Fig. 3 of Palmese and Conselice (2021)). The simulations of Ref. Khan and Holley-Bockelmann (2021) suggest that the formation and hardening of IMBH-IMBH pairs in the centers of dwarfs is very efficient and leads to their merger.
@@ II.1 Stellar consumption rate
Assuming most of the cluster’s mass is locked in light stars that survive for billions of years, the stellar number is $N_{\star}\simeq M_{\rm cl}/\overline{m}_{\star}$ where $\overline{m}_{\star}\simeq 0.6M_{\odot}$ for a Kroupa initial mass function Kroupa (2002). We neglect the contribution of BHs (and of the central BH, of mass $M_{\rm BH}$) to the total mass of the cluster, as the former can be shown to contribute no more than $\simeq 10\%$ of the total mass Kritos and Cholis (2020). According to the virial theorem, the root-mean-square velocity in the cluster is $v_{\rm rms}\simeq(0.4GN_{\star}\overline{m}_{\star}/r_{\rm h})^{1/2}$ [see, e.g., Eq. (2) from Ref. Spitzer and Hart (1971)]. We further assume isothermal and isotropic conditions within the half-mass radius, so that the velocity dispersion in the cluster core is $\sigma=v_{\rm rms}/\sqrt{3}$. The influence radius of the central BH is defined as $r_{\rm a}\equiv GM_{\rm BH}/\sigma^{2}$, and it corresponds approximately to the radius within which the potential of the BH starts to dominate over the stellar component. It also roughly corresponds to the distance from the center which encloses a total stellar mass of $\approx M_{\rm BH}$. Stars within $r_{\rm a}$ can have bound orbits around the central BH.
@@ II.2 Star cluster evolution
The tidally limited model can be applied to systems that experience the tidal field of the host galaxy within the inner kpc and inspiral to the center. On the other hand, the isolated model is more appropriate for the evolution of nuclear star clusters, since the tidal field cancels at the center of the galaxy. The isolated cluster evolution model is also adequate to describe clusters that reside several kpc away from the galaxy center; however, these isolated clusters in the halo of the galaxy are unlikely to merge with other clusters due to small coalescence cross sections.
@@
which can be integrated to get the time evolution of the stellar number
@@
Substituting (6) into (4) then yields $r_{\rm h}(t)$. When $\xi_{\rm e}=0$ and $\nu=0$ (i.e., in the case of no mass loss and no stellar evolution) this solution reduces to the late-time radius growth law $r_{\rm h}(t)\propto(t-\tau_{\rm cc})^{2/3}$: see, e.g., Eq. (4) in Antonini et al. (2019).
@@ II.4 Dependence on the cluster’s initial conditions
Figure 3: Contour plots of the BH mass assembled via runaway tidal encounters in the $(M_{\rm cl,0},r_{\rm h,0})$ plane at $t=0.1\tau_{\rm Hub}$ (left panels) and $t=\tau_{\rm Hub}$ (right panels). The top and bottom panels refer to isolated and tidally limited star clusters, respectively. The $\tau_{\rm BH}\geq\tau_{\rm Hub}$ ($\tau_{\rm BH}\geq 0.1\,\tau_{\rm Hub}$) region above the blue dashed line corresponds to clusters for which the time required for the BH to begin growing is larger than a Hubble time (one-tenth of a Hubble time) due to nonevaporation of the BH subsystem. In the top panels, the $\tau_{\rm cc}\leq 2\,\rm Myr$ region below the dashed orange line corresponds to clusters for which core collapse occurs before massive stars become BHs, and an IMBH might form through runaway stellar mergers instead. The $v_{\rm esc,0}>200\,\rm km\,s^{-1}$ region (dashed green line) corresponds to clusters for which the initial escape velocity is large enough for the repeated BH merger channel to dominate. In the bottom panels, the $\tau_{\rm ev}\leq\tau_{\rm Hub}$ ($\tau_{\rm ev}\leq 0.1\tau_{\rm Hub}$) region corresponds to clusters that evaporate on a timescale less than $\tau_{\rm Hub}$ ($0.1\tau_{\rm Hub}$): see Eq. (7). The parameter $f_{\rm s}$ is set to 0.5.
@@ II.5 Delay time
We consider the formation of IMBH-IMBH pairs through the coalescence of two star clusters, each cluster hosting one IMBH (see Fig. 1). This can be accomplished either through the merger of globular clusters that inspiral into the central regions of their galaxy through dynamical friction Ebisuzaki et al. (2001) or through the coalescence of nuclear star clusters in the postmerger phase of two dwarf galaxies Palmese and Conselice (2021). Alternatively, IMBHs may form in star clusters that fragment in the turbulent central regions of a galaxy and are then brought together to form a bound pair Shi et al. (2024).
@@
    If $E_{\rm b}=Gm_{1}m_{2}/(2a)$ is the binding energy of the IMBH-IMBH binary, then the rate of change of binding energy is given by $dE_{\rm b}/dt=\langle\Delta E_{\rm b}\rangle(dN/dt)$. Here the mean energy change per interaction is given by $\langle\Delta E_{\rm b}\rangle=(H/(4\pi))(m_{\star}/m_{12})E_{\rm b}$, where $H\simeq 15$ and $m_{12}=m_{1}+m_{2}$ Quinlan (1996). Moreover, $dN/dt$ is the binary-single interaction rate given by Eq. (6.15) of Ref. Merritt and Poon (2004), computed by setting the loss-cone radius equal to the semimajor axis of the binary, $a_{\rm B}$, and the central mass equal to the binary’s mass, $m_{12}$. Combining these equations leads to a differential equation for the evolution of $a_{\rm B}$
@@
    The semimajor axis shrinks until the emission of GWs is efficient enough to drive the evolution of the binary. At this point, the binary decouples from the dynamics, and we denote the corresponding semimajor axis by $a_{\rm gw}$. The eccentricity at decoupling, $e_{\rm gw}$, is sampled from a thermal distribution, and the hardening timescale $\tau_{\rm har}$ is computed as in Eq. (9) from Ref. Sesana and Khan (2015), where the density is set equal to $\rho_{\rm eff}$.
@@ III.1 Bayesian analysis
In this subsection we present the Bayesian framework used to infer the cluster properties. We start with a description of the cluster parameters (Sec. III.1.1), followed by the method we use to reweigh the prior (Sec. III.1.2) and a discussion of our chosen astrophysical prior (Sec. III.1.3).
@@ III.1.2 Source-parameter prior reweighing
When performing PE on our chosen events, we assume some prior on the source parameters, denoted by $p_{\rm PE}(\boldsymbol{\theta})$. Since the likelihood function remains invariant under the $\cal F$-map, so that $p(d|\boldsymbol{\theta})=p(d|\boldsymbol{\lambda})$, it follows that the priors on $\boldsymbol{\theta}$ and $\boldsymbol{\lambda}$ are related by a Jacobian as $p_{\rm PE}(\boldsymbol{\theta})d\boldsymbol{\theta}=p_{\rm PE}(\boldsymbol{\lambda})d\boldsymbol{\lambda}$, and therefore they are not independent from each other. This implies that a prior on $\boldsymbol{\lambda}$ is induced by the PE prior choice, which may not be consistent with the astrophysically motivated prior $p_{\rm astro}(\boldsymbol{\lambda})$ (that will be described in Sec. III.1.3 below). Hence, we reweigh the posterior samples on source parameters of our merger events by the astrophysical prior on $\boldsymbol{\theta}$, $p_{\rm astro}(\boldsymbol{\theta})$, which we introduce to be consistent with $p_{\rm astro}(\boldsymbol{\lambda})$. To obtain $p_{\rm astro}(\boldsymbol{\theta})$, we first draw samples $\{\boldsymbol{\lambda}_{i}\}$ from $p_{\rm astro}(\boldsymbol{\lambda})$, then we get the samples $\boldsymbol{\theta}_{i}={\cal F}(\boldsymbol{\lambda}_{i})$ with forward modeling, and finally we estimate the function $p_{\rm astro}(\boldsymbol{\theta})$ with kernel density estimation on these samples.
@@ IV.1.2 Runaway stellar collisions
A simple analytic model for the production of IMBHs through this channel, where the mass of the IMBH progenitor is given as a function of the cluster’s initial mass, was developed in Ref. Portegies Zwart and McMillan (2002) and verified through $N$-body simulations. Assuming that the very massive star assembled via this mechanism collapses directly into an IMBH, and neglecting any stellar wind effects that would reduce the mass of the star by the end of its life, Eq. (16) from Portegies Zwart and McMillan (2002) finds the simple relation $M_{\rm IMBH}\simeq m_{\rm seed}+0.008M_{\rm cl,0}$. Here $m_{\rm seed}$ is the zero-age main-sequence seed mass of the heaviest star that initially forms in the cluster, which may be $\sim 150M_{\odot}$ Figer (2005). Based on this equation, the mass scale of star clusters that contribute to the formation of hundred-$M_{\odot}$ IMBHs is similar to the mass scale of the runaway TDE channel examined in this paper for $f_{\rm s}=0.5$. Nevertheless, the stellar collision channel requires sufficiently small initial relaxation times (hence, very compact stellar systems), so that the process can kick-start before the heaviest stars form a BH subsystem that can prevent such a stellar runaway. The formation of the IMBH through stellar collisions is very rapid (within only a few Myr), and based on the classification of Ref. Giersz et al. (2015) it would correspond to a fast growth scenario.
@@ IV.1.4 Gas accretion
Observational data imply that, despite being dry stellar systems, globular clusters contain some gas in their centers, resulting from stellar winds of low-mass stars and ionized by ambient white dwarfs. For instance, Ref. Abbate et al. (2018) inferred a density of gas particles $n_{\rm e}=0.23\,\rm cm^{-3}$ using the dissipative effects that the gas would have on the emission of millisecond pulsars in 47 Tucanae.